# Advancing Multi-Party Dialogue Systems with Speaker-ware Contrastive Learning

### v1

@@ 4.3 Implementation Details
Our CMR model is implemented using PyTorch, with T5-base serving as the backbone. The learning rate is set to $4e^{-5}$ and we use the AdamW optimizer. The training procedure is divided into two stages. In Stage One, the encoder is trained with contrastive learning for 10 epochs with a batch size of 4. For each batch, 7 negative samples are used, with 4 samples coming from the same conversation and 3 from other samples within the same batch. After completing Stage One, Stage Two involves joint training with response generation and contrastive learning for 2 epochs, also with a batch size of 4. In this stage, 8 negative samples are used, distributed as follows: 1 sample from the previous utterance by the target speaker, 2 samples from other speakers within the same conversation, 3 samples from the same batch, and 2 model-generated beam search samples. The temperature $\tau$ is setting to 0.1. The model training was conducted on an RTX Titan GPU.
4.4 Automatic Evaluation Results
As shown in Table 1, to evaluate the performance of our proposed CMR model, we conducted experiments comparing it against several state-of-the-art models: BART, T5, CONT, HeterMPC, and MADNet. The evaluation metrics used include F1 Score, BLEU, and ROUGE-L, which provide a comprehensive assessment of the quality, relevance, and fluency of the generated responses. what is more, the results are averaged over five runs.
Compared to its backbone model, T5-base, CMR shows substantial improvements across all metrics, with some metrics nearly doubling. In particular, compared to CONT, which is also a contrastive learning-based response generation model, our model also shows significant improvements. Specifically, the F1 Score of CMR is 26.02% higher than that of CONT. For BLEU-1, BLEU-2, BLEU-3 and ROUGE-L, CMR achieves relative improvements of 39.26%, 27.16%, 31.13% and 20.58%, respectively. These results suggest that our approach to integrating contrastive learning with multi-party dialogue response generation is highly effective.
4.5 Human Evaluation Results
In addition to automatic evaluation, we conducted human evaluations (Table 2) to assess the quality of responses generated by our CMR model compared to baseline models. We used four metrics for evaluation: Relevance, Coherence, Engagement, and Speaker Appropriateness. Here, Speaker Appropriateness checks if the response matches the target speaker’s style and persona. Annotators rated each response on a scale of 1 to 5, with dialogue contexts specifically selected to require responses from one of the six main characters in "Friends." Annotators, blind to the model that generated each response and were instructed to reviewing training set utterances to familiarize themselves with speaking style of the main characters.
@@ 4.5 Human Evaluation Results
As shown in Table 2, our model outperformed the baseline models in all metrics. Notably, general pre-trained models like BART and T5 often produced dull responses, with around 65% of their outputs resembling generic phrases like "I am sorry" or "yes yes yes." This resulted in lower scores for relevance, engagement, and speaker appropriateness, despite relatively high coherence scores due to their logical consistency. Meanwhile, graph-based models like HeterMPC and MADNet struggled without additional annotations, with HeterMPC frequently generating meaningless phrases across all samples. Finally, while graph structures helped improve relevance and coherence, particularly for HeterMPC and MADNet, contrastive learning in models like CMR proved more effective in capturing different speaking styles, demonstrating the unique advantages of each approach.
@@ 5.1 Ablation Study
To understand the contribution of each component of our CMR model, we conducted a series of ablation experiments. Each ablation removes a specific component or modifies the model’s training process, allowing us to observe the impact on performance. The results of these ablation studies are summarized in Table 3.
@@ 5.1.2 Impact of Stage I
Removing the entire Stage I resulted in a significant performance drop across all metrics. The F1 score decreased by 0.97 points, BLEU-1 by 1.59 points, BLEU-2 by 0.38 points, BLEU-3 by 0.18 points, and ROUGE-L by 0.59 points. This demonstrates that the contrastive learning stage is crucial for improving the model’s understanding of dialogue context and speaker styles.
$N_{1},T_{1},N_{2},T_{2},P,B$ Metrics F1 BLEU-1 BLEU-2 BLEU-3 Rouge-L
$1,1,1,1,1,1$ 7.84 8.00 2.63 1.40 7.09
$2,2,2,1,1,1$ 7.95 8.47 2.69 1.37 7.20
$3,3,3,1,1,1$ 7.96 8.44 2.76 1.40 7.19
$\mathbf{4,4,4,2,1,2}$ 8.14 8.63 2.95 1.39 7.45
$5,5,5,3,1,3$ 8.17 8.64 2.87 1.41 7.51
$6,6,6,3,2,3$ 8.20 8.76 2.89 1.37 7.54
$8,8,8,4,3,4$ 8.02 8.58 2.77 1.34 7.20
@@ 5.2 Impact of the Number of Negative Samples
The number of negative samples is crucial for contrastive learning. Therefore, we explored the impact of varying the number of negative samples on the performance of CMR. As shown in Table 4, increasing the number of negative samples initially improves the experimental results. However, beyond a certain point, the performance begins to fluctuate and does not continue to improve steadily. This fluctuation may be due to the inclusion of more irrelevant or less similar negative samples as the number of negatives increases. For instance, as the value of k increases in beam search-generated top-k negative samples, more less optimal negative samples are included, which can negatively impact the effectiveness of contrastive learning.
@@ 5.3 Visualization of the Impact of Speaker aware in Stage I
To further understand the impact of Speaker aware in Stage I on the performance of the CMR model, we visualized the representations learned by the model with and without Stage I, as shown in Figure 3. Specifically, we visualized the attention weights of the $<$/s$>$ token to other tokens in the encoder. The reason for choosing the $<$/s$>$ token, as mentioned earlier, is based on our previous experiments. We found that during generation, the cross-attention weights of the decoder in the T5 model are primarily concentrated on the $<$/s$>$ token in the encoder. This indicates that the $<$/s$>$ token acts as an information aggregator in encoder, and its attention weights to other context tokens can effectively reflect the model’s primary focus. As shown in Figure 2, the attention weights of the $<$/s$>$ token in the CMR model with Stage I show a more concentrated pattern. The attention is focused on specific tokens that are contextually relevant, such as names (“Mona", “Ross") and pronouns (“I," “you"), indicating a better understanding of the dialogue context.
@@ 6 Conclusion
In this paper, we proposed CMR, a novel contrastive learning-based multi-party dialogue response generation model. Our approach leverages a two-stage training process to enhance the model’s understanding of dialogue context and speaker styles. In Stage One, we employ contrastive learning to differentiate between relevant and irrelevant utterances, significantly improving the model’s ability to comprehend complex multi-party dialogues. Stage Two combines response generation with contrastive learning, further refining the model’s capability to produce contextually relevant and coherent responses. In conclusion, our CMR model represents a significant advancement in multi-party dialogue response generation, offering a robust framework that effectively integrates contrastive learning to tackle the intricate dynamics of multi-party conversations. We believe that our findings will inspire further research and development in this challenging area, ultimately leading to more advanced and capable dialogue systems.
@@ 7 Limitations
  • The use of contrastive learning, particularly in a multi-stage training process, results in significantly increased training times compared to simpler models. Constructing negative samples and distributed training further adds to the computational cost.
  • Although the model performs well on the Friends dataset, which has clear and well-defined speaker identities, its ability to generalize to other domains or dialogues with less structured or noisier data has not been thoroughly explored.
  • While the model effectively captures speaker-specific characteristics in datasets with distinct characters, such as Friends, it may not perform as well in scenarios where speakers frequently change or where speaker-specific features are less defined. This limits its effectiveness in more spontaneous, open-domain dialogues.
@@ Appendix A The algorithm of entire training process
1:Stage One: Encoder Training
@@
15:Stage Two: Joint Training
@@
Appendix B Clean Friends Dataset
The Friends dataset we obtained included many low-quality samples. Some samples contained numerous meaningless and repetitive phrases, while others were too short and involved only two speakers. Additionally, some samples were fragmented and lacked context, making them unsuitable for training a robust dialogue model.
@@ Appendix B Clean Friends Dataset
To address these issues, we first set a list of “bad words," as shown in Table 5. If a sample’s response contained these bad words in more than 30% of its content, the sample was removed. Additionally, we removed samples where the dialogue involved only two speakers. Finally, we manually removed samples that lacked sufficient context.
Appendix C Configuration of Comparison Models
The original HeterMPC and MADNet models use BART-base as their backbone model. We first tested their performance with this configuration. Then, we replaced their backbone model with BART-large. During this modification, we adjusted the “num_gat_heads" parameter in their configuration files to 8, as the original setting of 6 was not divisible by the 1024-dimensional hidden state of BART-large. The results reported in the paper use BART-large as the backbone because both models achieved better performance with this configuration.
For the configurations of LLaMA-3 using Prompt and in-context learning, refer to Table 6. In our experiments, the prompt shown in the table yielded the best results and is used in the reported experiments. The in-context learning configuration involved testing with 1 to 6 samples, and we found that using 3 samples provided the best performance.

### v2

@@ 4.3 Implementation Details
Our CMR framework is combined with different backbone models, T5 and LLaMA 3.1. For LLaMA-based models, we apply LoRA during training to enable parameter-efficient fine-tuning. The learning rate is set to $4e^{-5}$ and we use the AdamW optimizer. In Stage I, the model is trained with contrastive learning for 10 epochs with a batch size of 4. For each batch, 7 negative samples are used, with 4 samples coming from the same conversation and 3 from other samples within the same batch. Stage II involves joint training with response generation and contrastive learning for 2 epochs, also with a batch size of 4 and $\lambda$ is setting to 2. In this stage, 8 negative samples are used, distributed as follows: 1 sample from the previous utterance by the target speaker, 2 samples from other speakers within the same conversation, 3 samples from the same batch, and 2 model-generated beam search samples. The temperature $\tau$ is set to 0.1.
4.4 Automatic Evaluation Results
As shown in Table 1, to evaluate the performance of our proposed CMR framework, we conduct experiments comparing it against several state-of-the-art models. The results are averaged over five runs.
On the Friends dataset, CMR demonstrates significant improvements across all evaluation metrics when combined with different backbone models, with several metrics nearly doubling. Specifically, compared to CONT, another contrastive learning-based dialogue response generation model, CMR-T5 substantially increases the F1 score by 26.02%, and achieves notable gains on BLEU and ROUGE scores. Additionally, compared to graph-based models such as HeterMPC and MADNet, CMR-T5 also exhibits substantial performance improvements. Meanwhile, the decoder-only LLM variant, CMR-LLaMA, significantly outperforms the vanilla LLaMA model, further highlighting the effectiveness and generalizability of our approach.
For the Ubuntu dataset, due to its data format being incompatible with the full CMR model, we apply only CMR (Stage II) for evaluation. Despite this limitation, CMR still outperform all other models of the same scale.
4.5 LLM-Based Pairwise Preference Evaluation
We observe that for such complex multi-party dialogue scenarios, conventional automatic evaluation metrics may be insufficient to fully capture the quality differences between model responses. To address this, we employ the GPT-4o model as an LLM-based judge, instructing it to carefully consider aspects such as speaker style consistency, personalization, contextual coherence, and response informativeness.
Specifically, we randomly sample 500 dialogue instances from the Friends dataset and conduct pairwise preference comparisons among responses generated by CMR-T5, CMR-LLaMA, LLaMA 3.1 8B, and LLaMA 3.1 70B. As shown in Figure 3, although CMR-LLaMA achieves only slightly better scores than CMR-T5 on automatic metrics, the LLM-based preference evaluation shows that CMR-LLaMA is preferred in 66% of the comparisons, indicating its superior ability to generate high-quality responses in multi-party dialogue settings.
4.6 Human Evaluation Results
In addition to automatic evaluation, we conduct human evaluations (Table 2) to assess the quality of responses generated by our CMR-T5 model compared to baseline models. We use four metrics for evaluation: Relevance, Coherence, Engagement, and Speaker Appropriateness. Here, Speaker Appropriateness checks if the response matches the target speaker’s style and persona. Annotators rate each response on a scale of 1 to 5, with dialogue contexts specifically selected to require responses from one of the six main characters in "Friends." Annotators, who are blind to the model that generated each response and were instructed to reviewing training set utterances to familiarize themselves with speaking style of the main characters.
Figure 3: LLM-based pairwise preference comparison on the Friends dataset.
@@ 4.6 Human Evaluation Results
As shown in Table 2, our model outperforms the baseline models in all metrics. Notably, general pre-trained models like BART and T5 often produce dull responses, with around 65% of their outputs resembling generic phrases like "I am sorry" or "yes yes yes." This results in lower scores for relevance, engagement, and speaker appropriateness, despite relatively high coherence scores due to their logical consistency. Meanwhile, graph-based models like HeterMPC and MADNet struggle without additional annotations, with HeterMPC frequently generating meaningless phrases across all samples. Finally, while graph structures help improve relevance and coherence, particularly for HeterMPC and MADNet, contrastive learning in models like CMR proves more effective in capturing different speaking styles, demonstrating the unique advantages of each approach.
@@ 5 Analysis and Discussion
All analyses in the following sections are conducted based on the CMR-T5 model.
@@ 5.1 Ablation Study
To understand the contribution of each component of our CMR model, we conduct a series of ablation experiments. Each ablation removes a specific component or modifies the model’s training process, allowing us to observe the impact on performance. As mentioned earlier, we conducted ablation experiments on Friends to isolate component impacts, as Ubuntu’s format limits full CMR application. The results of these ablation studies are summarized in Table 3.
@@ 5.1.2 Impact of Stage I
Removing the entire Stage I results in a significant performance drop across all metrics. The F1 score decreased by 0.97 points, BLEU-1 by 1.59 points, BLEU-2 by 0.38 points, BLEU-3 by 0.18 points, and ROUGE-L by 0.59 points. This demonstrates that the contrastive learning stage is crucial for improving the model’s understanding of dialogue context and speaker styles.
@@ 5.2 Impact of the Number of Negative Samples
The number of negative samples is crucial for contrastive learning. Therefore, we explored the impact of varying the number of negative samples on the performance of CMR. As shown in Appendix D.
@@ 5.3 Visualization of the Impact of Speaker aware in Stage I
To further understand the impact of Speaker aware in Stage I on the performance of the CMR-T5 model, we visualize the representations learned by the model with and without Stage I, as shown in Appendix E.
5.4 Generalization and Application Scenarios.
CMR is particularly well-suited for scenarios where a fixed set of speakers appears across dialogues, such as NPC conversations in games or episodic scripted content. In these cases, modeling consistent speaker styles across samples can bring notable benefits.
Nonetheless, CMR is not restricted to such settings. On the Ubuntu dataset where speakers rarely repeat, CMR still outperforms all baseline models of the same scale, despite being slightly behind LLM. We can say that the second stage of our framework alone is effective in modeling multi-party dialogues, even in the absence of speaker-level recurrence. Moreover, unlike prior methods, CMR requires no additional annotations, further demonstrating its adaptability to diverse dialogue scenarios.
To further evaluate the generalization ability of CMR, we conduct additional experiments in settings where the model encounters speakers with limited presence in the training data. As detailed in the appendix F, CMR still achieves strong performance under this condition. Intuitively, the phenomenon may be attributed to the model’s ability to generalize from frequently seen speakers. While the model may be unable to fully capture the speaking style of rarely seen speakers, it can effectively leverage its understanding of other speakers to distinguish between interlocutors. Thus, CMR remains capable of identifying speaker-specific factual content and maintaining consistency in speaker-related information.
5.5 Isolating Speaker Style: Comparison with Per-Speaker Models
One of the core motivations of CMR is to model individualized speaker style within multi-party dialogues. A natural question arises: would training a separate model for each speaker yield better style modeling, as each model can specialize in the language patterns of a single character?
To address this question, we conduct an experiment on the Friends dataset by training six separate models, one for each of the six main characters. During inference, we invoke the model corresponding to the ground-truth speaker. Please note that although the Friends dataset includes additional speakers beyond the six main ones, this experiment is restricted to training and testing only on the six main characters. As such, the results of CMR in this setting are slightly different from those reported earlier. The details are shown in Appendix G, the experimental results clearly demonstrate that training separate models for each speakers significantly poorer performance compared to CMR.
5.6 Robustness to Varying Speaker Counts
To evaluate the robustness of our model in handling dialogues with different levels of speaker complexity, we conducted additional experiments on subsets of dialogues containing 3, 4, 5, and 6 speakers.
Using the Entire Dataset.
We first tested our model, trained on the entire Friends dataset, on conversations with varying speaker counts. We observe a gradual performance decrease as the number of speakers increases. However, the decline may be attributed not only to the inherent complexity of multi-party dialogue but also to the training data distribution, which predominantly features conversations with 3 or 4 speakers.
Using Balanced Subsets.
To isolate the effect of speaker count, we created balanced subsets of 400 samples for each speaker setting. We retrained the model on these balanced subsets. Although the overall scores are lower compared to the model trained on the full dataset, the performance decline across speaker counts is notably smaller, suggesting that our model retains a reasonable degree of robustness when faced with increasing dialogue complexity.
All details and results are shown in Appendix H.
5.7 Training Cost and Deployment
Due to CMR’s two-stage training strategy and the need to generate responses for contrastive learning during the second stage, CMR requires more training time than its backbone model.
However, once training is complete, CMR demonstrates a distinct advantage over graph-based models during deployment. The contrastive learning component of CMR is entirely absent during inference, meaning that its resource requirements and inference speed are identical to those of a standard T5 model. In contrast, graph-based models require the construction of graphs during both training and inference, which incurs additional computational overhead.
@@ 6 Conclusion
We propose CMR, a novel contrastive learning-based multi-party dialogue response generation framework. Our approach leverages a two-stage training process to enhance the model’s understanding of dialogue context and speaker styles. In Stage I, we employ contrastive learning to differentiate between relevant and irrelevant utterances, significantly improving the model’s ability to comprehend complex multi-party dialogues. Stage II combines response generation with contrastive learning, further refining the model’s capability to produce contextually relevant and coherent responses. In conclusion. Our CMR framework advances multi-party dialogue response generation by integrating contrastive learning to effectively handle complex dialogue dynamics and speaker variability, providing a robust foundation for future research in this domain.
@@ 7 Limitations
  • While the model does not require additional resources compared to the backbone model during deployment, its training process is time-consuming due to the multi-stage contrastive learning framework.
  • While the model effectively captures speaker-specific characteristics in datasets with distinct characters, such as Friends, and also demonstrates promising results on Ubuntu by leveraging only Stage II training, it may still be inherently better suited for scenarios where the set of speakers is relatively fixed and speaker-specific features are explicitly available.
@@ Appendix A The algorithm of entire training process
1:Stage I: Encoder Training
@@
15:Stage II: Joint Training
@@
Appendix B Clean Dataset
The Friends and Ubuntu dataset we obtained included many low-quality samples. Some samples contained numerous meaningless and repetitive phrases, while others were too short and involved only two speakers. Additionally, some samples were fragmented and lacked context, making them unsuitable for training a robust dialogue model.
@@ Appendix B Clean Dataset
$N_{1},T_{1},N_{2},T_{2},P,B$ Metrics F1 BLEU-1 BLEU-2 BLEU-3 Rouge-L
$1,1,1,1,1,1$ 7.84 8.00 2.63 1.40 7.09
$2,2,2,1,1,1$ 7.95 8.47 2.69 1.37 7.20
$3,3,3,1,1,1$ 7.96 8.44 2.76 1.40 7.19
$\mathbf{4,4,4,2,1,2}$ 8.14 8.63 2.95 1.39 7.45
$5,5,5,3,1,3$ 8.17 8.64 2.87 1.41 7.51
$6,6,6,3,2,3$ 8.20 8.76 2.89 1.37 7.54
$8,8,8,4,3,4$ 8.02 8.58 2.77 1.34 7.20
@@
To address these issues, we first set a list of “bad words," as shown in Table 4. If a sample’s response contained these bad words in more than 30% of its content, the sample was removed. Additionally, we removed samples where the dialogue involved only two speakers. Finally, we manually removed samples that lacked sufficient context. After data pre-processing, there are 5000 conversations for training and 653 conversations for testing on Friends dataset. For the Ubuntu dataset, since each utterance is relatively short, we also require that each dialogue contains at least 6 utterances and the entire conversation needed to exceed 512 tokens. There are about 67000 conversations for Ubuntu dataset.
Appendix C Configuration of LLaMA
We use LLaMA 3.1 as our comparison model. For the configurations of LLaMA-3 using Prompt and in-context learning, refer to Table 6. In our experiments, the prompt shown in the table yielded the best results and is used in the reported experiments. The in-context learning configuration involved testing with 1 to 6 samples, and we found that using 3 samples provided the best performance.
Appendix D Impact of the Number of Negative Samples
As shown in Table 5, increasing the number of negative samples initially improves the experimental results. However, beyond a certain point, the performance begins to fluctuate and does not continue to improve steadily. This fluctuation may be due to the inclusion of more irrelevant or less similar negative samples as the number of negatives increases. For instance, as the value of k increases in beam search-generated top-k negative samples, more less optimal negative samples are included, which can negatively impact the effectiveness of contrastive learning.
Appendix E Visualization of the Impact of Speaker aware in Stage I
As shown in Figure 4, we visualize the attention weights of the $<$/s$>$ token to other tokens in the encoder. The reason for choosing the $<$/s$>$ token, as mentioned earlier, is based on our previous experiments. We found that during generation, the cross-attention weights of the decoder in the T5 model are primarily concentrated on the $<$/s$>$ token in the encoder. This indicates that the $<$/s$>$ token acts as an information aggregator in encoder, and its attention weights to other context tokens can effectively reflect the model’s primary focus. As shown in Figure 2, the attention weights of the $<$/s$>$ token in the CMR model with Stage I show a more concentrated pattern. The attention is focused on specific tokens that are contextually relevant, such as names (“Mona", “Ross") and pronouns (“I," “you"), indicating a better understanding of the dialogue context.
Appendix F Generalization to Rare and Unseen Speakers
To investigate the generalization capabilities of CMR with respect to rarely observed or completely unseen speakers during training, we extract dialogues from the Friends dataset involving speakers beyond the six main characters. Specifically, we isolate response samples from the test set where the speakers are infrequently encountered or entirely absent in the training phase. These samples are evaluated using the fully trained CMR model without any further fine-tuning and the results in Table 7 illustrate the performance of CMR on responses by rare or unseen speakers.
Evaluation F1 BLEU-1 BLEU-2 BLEU-3 ROUGE-L
Rare Speakers 8.11 8.65 2.85 1.38 7.29
Standard 8.14 8.63 2.95 1.39 7.45
Table 7: CMR Performance on responses from rare and unseen speakers.
The experimental outcomes demonstrate that the CMR model maintains strong performance in generating responses even for speakers rarely or never encountered during training.
A plausible explanation for this generalization ability is that although the CMR model may not fully capture the speaking style of rarely seen or unseen speakers due to limited data exposure, it can effectively leverage the comprehensive understanding gained from frequent speakers. With the assistance of the first stage, while the model may be unable to fully capture the speaking style of rarely seen speakers, it can effectively leverage its understanding of other speakers to distinguish between interlocutors. The second stage, is less affected by unseen speakers because it primarily focuses on dialogue themes, context continuity, and factual consistency. Consequently, CMR remains capable of recognizing thematic shifts and accurately maintaining speaker-related factual content, thereby ensuring coherent and contextually relevant responses despite encountering rarely or never-before-seen speakers.
Appendix G Training separate models for each speakers
We extract samples for the six main characters from the Friends dataset and train six baseline models separately. During the testing phase, a different baseline model will be invoked according to the speaker to evaluate its performance.
Method F1 Score BLEU-1 BLEU-2 BLEU-3 ROUGE-L
Separate models 7.33 7.90 2.61 1.25 7.04
CMR 8.14 8.63 2.95 1.39 7.45
Table 8: Comparison between separate models and our proposed method.
The experimental results clearly demonstrate significantly poorer performance compared to CMR. We attribute this outcome to two primary factors:
  • The approach of training separate models for each character only enables the model to capture superficial personalized information about individual characters. More importantly, it fails to help the model understand the nuanced differences between different characters.
  • The original Friends dataset has a limited number of samples to begin with. After splitting the training process, each model only has access to approximately one-sixth of the original data.
Appendix H Robustness to Varying Speaker Counts
In this section, we provide additional details and full results for our robustness analysis on varying speaker counts.
Experimental Setup.
We conducted two sets of experiments to assess the model’s performance when handling conversations with different numbers of participants (3, 4, 5, and 6 speakers):
  • Experiment 1: Using the Entire Dataset. We evaluated the model trained on the full dataset without controlling for speaker count distribution. This reflects a realistic but potentially biased setting, as most training samples involve only 3 or 4 speakers.
  • Experiment 2: Using Balanced Subsets. To control for data imbalance, we created balanced subsets with 400 samples each for conversations involving 3, 4, 5, and 6 speakers. We retrained the model on each subset and evaluated its performance.
Results on the Entire Dataset.
# Speakers F1 BLEU-1 BLEU-2 BLEU-3 ROUGE-L
3 8.50 9.00 3.10 1.50 7.80
4 8.21 9.13 2.99 1.47 7.49
5 7.91 8.47 2.88 1.35 7.36
6 7.65 8.22 2.73 1.30 7.10
Table 9: Performance on conversations with 3–6 speakers using the entire dataset.
Results on Balanced Subsets.
# Speakers F1 BLEU-1 BLEU-2 BLEU-3 ROUGE-L
3 7.84 8.46 2.93 1.48 7.28
4 7.69 8.23 2.82 1.41 7.12
5 7.69 8.13 2.77 1.39 7.09
6 7.63 8.22 2.73 1.37 7.11
Table 10: Performance on balanced subsets with 3–6 speakers (400 samples each).
Analysis.
The results across both experimental settings reveal a consistent performance drop as the number of speakers increases. However, when using balanced subsets, the performance degradation is much less pronounced, suggesting that our model can generalize reasonably well across varying